### v3

# VideoMark: A Distortion-Free Robust Watermarking Framework for Video Diffusion Models

Abstract
This work introduces VideoMark, a distortion-free robust watermarking framework for video diffusion models. As diffusion models excel in generating realistic videos, reliable content attribution is increasingly critical. However, existing video watermarking methods often introduce distortion by altering the initial distribution of diffusion variables and are vulnerable to temporal attacks, such as frame deletion, due to variable video lengths. VideoMark addresses these challenges by employing a pure pseudorandom initialization to embed watermarks, avoiding distortion while ensuring uniform noise distribution in the latent space to preserve generation quality. To enhance robustness, we adopt a frame-wise watermarking strategy with pseudorandom error correction (PRC) codes, using a fixed watermark sequence with randomly selected starting indices for each video. For watermark extraction, we propose a Temporal Matching Module (TMM) that leverages edit distance to align decoded messages with the original watermark sequence, ensuring resilience against temporal attacks. Experimental results show that VideoMark achieves higher decoding accuracy than existing methods while maintaining video quality comparable to watermark-free generation. The watermark remains imperceptible to attackers without the secret key, offering superior invisibility compared to other frameworks. VideoMark provides a practical, training-free solution for content attribution in diffusion-based video generation. Our code and data are available at VideoMark.
1 Introduction
In recent years, diffusion models have revolutionized the landscape of AI-generated content, emerging as the state-of-the-art technology for image and video generation [ho2020denoising, ho2022video, sohl2015deep, liu2025javisgpt]. These models can create highly realistic content that is increasingly indistinguishable from human-created media [rombach2022high]. The rapid advancement in generation quality has created an urgent need to track and attribute AI-generated content, particularly given growing concerns about copyright infringement and potential misuse [almutairi2022review, zhang2025cohemark]. To address these challenges, watermarking techniques have emerged as a crucial solution for ensuring content traceability and authentication in the era of AI-generated media.
Figure 1: VideoMark outperforms VideoShield across three key metrics: message length, robustness, and invisibility.
Traditional watermarking methods for both images and videos typically operate as post-processing techniques, where watermarks are embedded after content generation [luo2023dvmark, zhang2019robust, ye5194026cmat]. These methods not only require additional computational overhead but also suffer from limited generalization capabilities. Recent research has shifted towards embedding watermarks during the generation process itself. Leveraging the reversibility of DDIM [song2020denoising], several methods have achieved success in the image domain by manipulating the initial Gaussian noise—e.g., Tree-Ring [wen2023tree]—or embedding messages into the noise distribution, as in Gaussian Shading [yang2024gaussian] and PRC-Watermark [gunn2024undetectable].
However, directly adapting image watermarking techniques to the video domain presents unique challenges. First, video DDIM inversion yields lower accuracy than image-based methods. And as a result, methods like VideoShield [hu2025videoshield] repeat watermark patterns in initial noise to enhance detection, but these compromise video quality and watermark invisibility. Second, watermark robustness suffers against temporal attacks like frame deletion or reordering, because treating videos as a single entity fails to localize watermarks temporally. Third, variable video lengths pose difficulties for algorithms relying on fixed noise initialization, limiting scalability.
To address video watermarking challenges, we first define essential characteristics for our proposed watermark. Primarily, the watermark embedded within the initial latent noise must cause negligible perturbation to the original noise space. Secondly, our approach involves inserting unique watermarks into individual frames. Beyond this per-frame embedding, we must also establish a temporal relationship for these watermarks across consecutive frames to improve robustness.
With these needs in mind, we introduce VideoMark, a distortion-free robust watermarking framework designed for video diffusion models. To achieve an imperceptible watermark that preserves the original noise characteristics, VideoMark utilizes pseudorandom error correction (PRC) codes [christ2024pseudorandom]. These codes map the watermark bits directly onto the initialized Gaussian noise for every frame. This specific design ensures the watermark integrates seamlessly, thus fulfilling our first design goal. To enable frame-specific watermarking, VideoMark processes each frame’s watermark independently while preserving sequential consistency across frames. Specifically, we generate an extended watermark message sequence. For each video, a random starting position within this master sequence initializes the first frame’s watermark, and subsequent frames derive their watermarks sequentially. This aligns with our second design objective, facilitating both individualized frame watermarking and temporal coherence.
To accurately extract the watermark, we propose a temporal matching module (TMM), which uses edit distance to align the decoded message with the embedded watermark sequence, thereby improving decoding accuracy. Even under temporal attacks such as frame deletion, TMM preserves the robustness of the embedded watermark.
In our experiments, we evaluate the effectiveness of our watermarking framework across different video diffusion models, demonstrating high decoding accuracy, high-quality generated videos, and strong invisibility. Our watermark achieves higher decoding accuracy compared to VideoShield, which is currently the state-of-the-art watermarking approach for video diffusion models. Additionally, our watermark achieves the best video quality on both the objective video evaluation benchmark VBench[huang2024vbench] and subjective assessments, maintaining parity with watermark-free videos. Importantly, our watermark remains undetectable to attackers without the key, ensuring stronger imperceptibility than other watermarking frameworks.
In summary, the contributions of this work are summerized as follows:
  • We propose VideoMark, which leverages pseudo-random Gaussian space initialization to achieve undetectable watermarking in video diffusion models.
  • We introduce a frame-wise watermarking strategy with extended message sequences, solving the challenge of variable-length videos and temporal attacks.
  • Our extensive experiments demonstrate that VideoMark achieves higher decoding accuracy than existing methods while maintaining video quality on par with watermark-free generation across various video diffusion models and attack scenarios.
2 Related Work
2.1 Video Diffusion Models
Diffusion models [sohl2015deep] progressively add noise to map data distributions to a Gaussian prior and recover original data via iterative denoising. Video diffusion models [ho2022video] use a 3D U-Net with interleaved spatial and temporal attention to generate high-quality, temporally consistent frames. Building on latent diffusion models [rombach2022high], SVD [blattmann2023stable] learns a multi-dimensional latent space for high-resolution frame synthesis. During generation, DDIM sampling [song2021denoising] efficiently reduces sampling steps while maintaining video quality compared to DDPM sampling [ho2020denoising].
Video diffusion models primarily follow two paradigms: Text-to-Video[wang2023modelscope, hu2025videoshield, huang2024vbench], where videos are generated based on text prompts, and Image-to-Video[2023i2vgenxl, hu2025videoshield, blattmann2023stable], where a video is generated starting from a single image. These paradigms enable the generation of realistic videos, but they also raise concerns regarding the potential generation of misleading content or copyright infringement.
2.2 Video Watermark
Video watermarking technology embeds imperceptible patterns into visual content and employs specialized detection methods to verify watermark presence [liu2024survey]. These methods are typically classified into two paradigms: post-processing and in-processing schemes.
Post-processing schemes introduce minimal visual perturbations, typically at the pixel level. Recent works [videoseal, luo2023dvmark, zhang2019robust, revmark] focus on training watermark embedding networks by optimizing discrepancies between watermarked and original videos, as well as encoding-decoding differences. However, these methods may struggle to balance trade-offs between video quality and watermark robustness.
In contrast to post-processing methods, in-processing schemes integrate the watermarking into the video generation process of current generative video models to better utilize their capabilities. For instance, Videoshield[hu2025videoshield] extends the Gaussian Shading technique[yang2024gaussian] from the image domain to the video domain, achieving improved robustness. However, repeating watermark bits during initialization induces fixed latent patterns, consequently degrading the quality of generated videos. Currently, only one in-processing video watermarking method exists, and prior approaches struggle to balance watermark robustness with video quality. We propose an undetectable video watermarking scheme to address this trade-off.
3 Preliminaries
3.1 Diffusion Models
Diffusion models generate content through an iterative denoising process. Given a noise schedule ${\beta_{t}}{t=1}^{T}$, the forward process gradually adds noise to data $\mathbf{x_{0}}$:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$ (1)
The reverse process is learned to gradually denoise from $\mathbf{x}T\sim\mathcal{N}(0,\mathbf{I})$ to generate content. While DDPM [ho2020denoising] introduces stochasticity in each denoising step, DDIM [song2020denoising] provides an approximately invertible deterministic sampling process:
$\begin{split}\mathbf{x}_{t-1}=\ &\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}\right)\\ &+\sqrt{1-\alpha_{t-1}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\end{split}$ (2)
This deterministic reversibility enables control over the generation process through manipulation of the initial noise.
3.2 Pseudorandom Codes (PRC)
A PRC is a coding scheme that maps messages to statistically random-looking codewords. We adopt the construction from christ2024pseudorandom, which provides security based on the hardness of the Learning Parity with Noise (LPN) problem.
The PRC framework consists of three core algorithms:
  • $\text{KeyGen}(n,m,\text{fpr},t)\rightarrow\text{key}$: Generates a key for encoding $m$-bit messages into $n$-bit codewords with sparsity parameter $t$
  • $\text{Encode}(\text{key},\mathbf{m})\rightarrow\mathbf{c}$: Maps message $\mathbf{m}$ to codeword $\mathbf{c}\in\{-1,1\}^{n}$
  • $\text{Decode}(\text{key},\mathbf{s})\rightarrow\mathbf{m}$ or $\emptyset$: Recovers message from potentially corrupted signal $\mathbf{s}\in[-1,1]^{n}$
Our implementation supports soft decisions on recovered bits, optimized for robust watermarking (see supplementary materials for details).
3.3 Generative Video Watermarking and Threat Model
Diffusion-based video watermarking involves three key functions in the watermarking process:
1. Generation: $V=\mathcal{G}(m,k)$, where $\mathcal{G}$ generates a watermarked video $V$ by embedding message $m$ using secret key $k$ during the diffusion process.
2. Decoding: $\hat{m}=\mathcal{D}_{PRC}(V)$, where $\mathcal{D}_{PRC}$ extracts the watermark message $\hat{m}$ from the given video $V$.
[…]
We consider active adversaries who may perform various modifications on the watermarked video to remove or corrupt the embedded watermark. These include:
  • Temporal Attacks: Frame drop, insert, or swap, which disrupts the temporal structure.
  • Spatial Attacks: Frame-wise manipulations such as Gaussian blurring, colour jittering, and resolution compression, which aim to distort the watermark signal by degrading the visual content of individual frames.
Our framework aims to be robust against these attacks while ensuring the watermark remains imperceptible and the video quality is preserved.
4 Proposed Method
Figure 2: The overall framework of VideoMark. During the watermark embedding phase, $\epsilon$ denotes the standard Gaussian noise sampled randomly. In the I2V task, the first video frame prompts the prediction of initial noise during watermark extraction.
In this section, we provide a detailed explanation of the proposed unbiased watermarking method in video diffusion models. Specifically, in Section 4.1, we detail the process the watermark generation. In Section 4.2, we introduce the watermark extraction process.
4.1 Watermark Generation
Algorithm 1 Watermarked Video Generation
1:PRC-key $k$, number of frames $f$, channels $c$, height $h$, width $w$, message $m$, diffusion model $\mathcal{M}$, VAE decoder $\mathcal{D}$
2:Watermarked video $V$
3:Generate extended message sequence $M$ longer than maximum supported length
4:Randomly select starting position $p$ in $M$
5:for $i=1$ to $f$ do
6:  Extract frame message $m_{i}$ from $M$ starting at position $p+i$
7:  Encode $m_{i}$ into PRC codeword $c_{i}\in\{-1,1\}^{c\times h\times w}$
8:  Sample $\epsilon_{i}\sim\mathcal{N}(0,1)\in\mathbb{R}^{c\times h\times w}$
9:  Compute $\hat{\epsilon}_{i}\leftarrow c_{i}\cdot|\epsilon_{i}|$
10:end for
11:Denoise $\hat{\epsilon}$ using diffusion model $\mathcal{M}$ and decode with VAE decoder $\mathcal{D}$ to obtain video $V$
12:return $V$
In this section, we introduce the watermark generation process. VideoMark achieves high invisibility and video quality in diffusion-based watermarking by initializing each frame with pseudo-random Gaussian noise using PRC, followed by DDIM denoising[song2020denoising] and VAE decoding [vae]. To enhance diversity and adapt to varying video lengths, we employ an extended message list with a random start index.
Prior watermarking methods (e.g. VideoShield [hu2025videoshield]) often repeat identical noise patterns, compromising pseudo-randomness, reducing watermark bit capacity, and degrading invisibility and video quality. VideoMark addresses this by generating frame-specific pseudo-random initializations. For a video with $f$ frames, dimensions $c\times h\times w$ (channels, height, width), and a message bit $m_{i}^{\prime}$ per frame, the process is as follows. For each frame $i\in\{1,\ldots,f\}$, we sample Gaussian noise $\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(0,\mathbf{I})\in\mathbb{R}^{c\times h\times w}$. Using a PRC key $k$, we encode $m_{i}^{\prime}$ to obtain a codeword $\mathbf{c}_{i}=\text{Encode}(k,m_{i}^{\prime})\in\{-1,1\}^{c\times h\times w}$. The watermarked noise is:
$\hat{\boldsymbol{\epsilon}}_{i}=\mathbf{c}_{i}\cdot|\boldsymbol{\epsilon}_{i}|,$
where $\mathbf{c}_{i}$ modulates the sign of $\boldsymbol{\epsilon}_{i}$, preserving its magnitude. The noise sequence $\hat{\boldsymbol{\epsilon}}=[\hat{\boldsymbol{\epsilon}}_{1},\ldots,\hat{\boldsymbol{\epsilon}}_{f}]$ is denoised using a DDIM diffusion model $\mathcal{M}$. For each frame, DDIM iterates over $T$ steps:
$\begin{split}\hat{\boldsymbol{z}}_{i}^{(t-1)}=\ &\sqrt{\alpha_{t-1}}\left(\frac{\hat{\boldsymbol{z}}_{i}^{(t)}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{i}^{(t)},t)}{\sqrt{\alpha_{t}}}\right)\\ &+\sqrt{1-\alpha_{t-1}}\,\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{i}^{(t)},t),\end{split}$ (3)
with $\hat{\boldsymbol{z}}_{i}^{(T)}=\hat{\boldsymbol{\epsilon}}_{i}$, producing latent $\hat{\boldsymbol{z}}_{i}^{(0)}$. The VAE decoder $\mathcal{D}$ generates the watermarked video $V=[\mathcal{D}(\hat{\boldsymbol{z}}_{1}^{(0)}),\ldots,\mathcal{D}(\hat{\boldsymbol{z}}_{f}^{(0)})]$.
To adapt to videos of varying lengths and increase diversity, we generate an extended message list $M=[m_{1},\ldots,m_{L}]$, where $L>f_{\text{max}}$ and $f_{\text{max}}$ is the maximum supported frame count. For each video, we sample a start index $p\sim\text{Uniform}(0,L-f)$, selecting messages $m_{i}^{\prime}=M[p+i]$ for $i\in\{1,\ldots,f\}$. These are encoded via PRC to produce $\mathbf{c}_{i}$. The random start index ensures diverse initializations across videos, improving security and reducing detectable patterns, while supporting arbitrary video lengths. This frame-wise approach resists temporal and spatial attacks. The pipeline is shown in Figure 2 and Algorithm 1.
[…]
Table 1: Main results of VideoMark. All columns present bit accuracy metrics except the Video Quality column.
4.2 Watermark Extraction
In this section, we present our watermark extraction process which consists of three key functions: decoding, detection, and recovery. This approach effectively handles various attacks that may disrupt the video structure.
Decoding Function $\hat{m}=\mathcal{D}_{PRC}(V)$ extracts the embedded message from a watermarked video $V$ with $f$ frames. We first recover the approximate initial noise for each frame using the DDIM inverse process:
$\tilde{\boldsymbol{\epsilon}}_{i}=\mathcal{M}^{-1}(V_{i}),\quad i\in\{1,\ldots,f\}.$ (4)
Then, we decode each frame’s message bit using the sign pattern of the recovered noise:
$\hat{m}_{i}=\text{PRC.Decode}(k,\text{sign}(\tilde{\boldsymbol{\epsilon}}_{i}))$ (5)
where the Decode function extracts the message bit encoded in the sign pattern using the PRC key $k$ (details of the PRC algorithm can be found in supplementary materials), matching the encoding process where $\hat{\boldsymbol{\epsilon}}_{i}=\mathbf{c}_{i}\cdot|\boldsymbol{\epsilon}_{i}|$ and $\mathbf{c}_{i}\in\{-1,1\}^{c\times h\times w}$. The complete decoded sequence is returned as $\hat{m}=[\hat{m}_{1},\ldots,\hat{m}_{f}]$.
[…]
$p=\text{rank}(d_{\text{edit}}(m,\hat{m}))/N$ (6)
[…]
The edit distance calculation incorporates frame-wise Hamming distance, defined as:
$d_{H}(m_{i},\hat{m}_{j})=\frac{1}{|m_{i}|}\sum_{k=1}^{|m_{i}|}\mathbf{1}_{m_{i}[k]\neq\hat{m}_{j}[k]}$ (7)
This distance is normalized through a continuous mapping:
$d_{N}(m_{i},\hat{m}_{j})=2\big(d_{H}(m_{i},\hat{m}_{j})-0.5\big)$ (8)
Since two random binary sequences are expected to have a Hamming distance of 0.5, this transformation linearly scales distances to the range [1,1][-1,1], enhancing the sensitivity of the detection mechanism.
[…]
$I_{j}=\arg\min_{j}\{d_{\text{edit}}(m,\hat{m}_{j})\},\quad j\in\{1,\ldots,f\}$ (9)
Then, using the indices, we identify both the starting index $s$ and the optimal alignment path between $m$ and $\hat{m}$:
[…]
where $\mathcal{P}$ represents the sequence of operations (match, insert, delete, substitute) that transforms $m[s:]$ into $\hat{m}$ with minimal cost. Using this path, we recover the original message by extracting the corresponding subsequence from $m$ that aligns with $\hat{m}$:
$m^{\prime}=\{m[s+j]\mid\mathcal{P}_{j}\text{ is a match or substitute operation}\}$ (11)
This extracts precisely the elements from the original message that correspond to the decoded sequence after accounting for any frame manipulations.
5 Experiments
5.1 Experimental Setting
Implementation details. In our primary experiments, we explore both text-to-video (T2V) and image-to-video (I2V) generation tasks, employing ModelScope (MS) [wang2023modelscope] for T2V synthesis and I2VGen-XL [2023i2vgenxl] for I2V generation. The generated videos consist of 16 frames, each with a resolution of 512 $\times$ 512. The inference and inversion steps are set to their default values of 25 and 50, respectively. Watermarks of 512 bits are embedded into each generated frame of the two models. As described in the Section 4.2, we leverage DDIM inversion to obtain predicted initial noise. The threshold $\tau$ is set to 0.005. The number of random sequences $N$ is set to 1000. All experiments are conducted on an NVIDIA Tesla A800 80G GPU.
Baseline. We selected four watermarking methods as baselines for comparison: RivaGAN[zhang2019robust], REVMark[revmark], VideoSeal[videoseal], and VideoShield[hu2025videoshield]. All selected methods are open-source and specifically designed to embed multi-bit strings within a video. Specifically, we set 32 bits for RivaGAN, 96 bits for REVMark, 96 bits for VideoSeal and 512 bits for VideoShield. Among these methods, VideoShield is the only in-generation approach, whereas the others are post-processing techniques that necessitate training new models for watermark embedding.
Datasets. We select 50 prompts from the test set of VBench[huang2024vbench], covering five categories: Animal, Human, Plant, Scenery, and Vehicles, with 10 prompts per category. For the T2V task, we generate four videos for each prompt for evaluation, ensuring diversity in outputs while maintaining consistency in prompt interpretation. For the T2V task, we first leverage a text-to-image model Stable Diffusion 2.1[Rombach_2022_CVPR], to generate images corresponding to the selected prompts. These generated images are subsequently utilized to create videos. Overall, we generate a total of 200 videos for both tasks for the primary experiments. Additionally, for each prompt category in VBench, we generate 10 watermarked and 10 non-watermarked videos, resulting in a total of 8,000 watermarked and 8,000 non-watermarked videos for the watermark learnability comparison experiment.
Metric. We leverage Bit Accuracy to evaluate the ratio of correctly extracted watermark bits. To evaluate the quality of the generated videos, we conducted both objective and subjective assessments. For the objective evaluation, we leverage the metrics Subject Consistency, Background Consistency, Motion Smoothness, and Image Quality from VBench (see supplementary materials for details). For the subjective evaluation, we meticulously designed a pipeline that leverages GPT-4o to evaluate and score the generated videos (see supplementary materials for details).
5.2 Main Results
In Table 1, we present the main experimental results of VideoMark, including extraction accuracy, video quality, and both temporal and spatial robustness.
[…]
Table 2: GPT-4o-based quality assessment using an LLM-as-a-judge setting. Models: RG (RivaGAN), RM (REVMark), VS (VideoSeal), VSh (VideoShield), and VM(VideoMark).
[…]
Quality. The “Video Quality” columns present the objective experimental results of various watermarking methods on the VBench benchmark. VideoMark consistently achieves state-of-the-art performance across all four metrics in both tasks. In the I2V task, it surpasses the best post-processing method, VideoSeal, by 0.004, and outperforms the leading in-processing method, VideoShield, by 0.036. Notably, in terms of Image Quality (IQ), our method achieves scores of 0.692 on T2V and 0.581 on I2V.
In addition to the objective evaluation metrics, we adopt an LLM-as-a-judge strategy for subjective video quality assessment. From the 8,000 videos generated by each model, we randomly sample 1,000 videos and leverage GPT-4o evaluate their perceptual quality. We present the number of samples for which each method receives the highest score in Table 2. The results show that VideoMark achieves the most top-rated samples in both tasks, with 288 and 245 samples, respectively 48 more than the second-best method, VideoShield, in the T2V task, and 33 more than the second-best method, RivaGAN, in the I2V task. The visual results are provided in supplementary material.
Robustness. The “Temporal Tampering” and “Spatial Tampering” columns show robustness results under temporal and spatial attacks, respectively (detailed experimental settings are in supplementary materials).
Table 3: VideoMark robustness under temporal tampering attacks, reported as $p$-values.
ModelScope I2VGen-XL
Swap Insert Drop Swap Insert Drop
0.001 0.001 0.001 0.001 0.001 0.001
Table 4: Comparison of temporal robustness between VideoShield and VideoMark, using matching accuracy as the evaluation metric.
Method ModelScope I2VGen-XL
Swap Insert Drop Swap Insert Drop
VideoShield 1.000 1.000 1.000 0.983 0.983 0.981
[…]
For temporal tampering, we show the bit accuracy between the decoded and embedded message. As REVMark does not release the necessary model files, and VideoShield cannot handle videos with variable frames during decoding, we omit their results from this evaluation. The results show that VideoMark maintains a perfect bit accuracy of 1.000 in the T2V task. In the I2V task, it achieves an average bit accuracy of 0.996 and retains a strong performance of 0.991 even under the most challenging frame insertion attack. These findings further suggest that Videomark can reliably decode the embedded message even under temporal tampering, thereby ensuring that the watermark is robustly distributed across frames.
In addition, in Table 3, we present the p-values of VideoMark’s detection results under temporal tampering attacks. Both models exhibit a p-value of 0.001 in detecting temporal tampering, which indicates strong statistical significance. Table 4 compares frame matching accuracy between VideoMark and VideoShield. Results show VideoMark achieves up to 0.996 accuracy in the I2V task, demonstrating the TMM module’s effectiveness in reliably reconstructing the original temporal order.
For spatial tampering (details in supplementary materials), VideoMark embeds 32 bits per frame, achieving perfect bit accuracy (1.000) on the T2V task. In the I2V task, despite a lower score under Gaussian Blur (0.857), it attains the highest average accuracy (0.911) across all attacks.
Invisibility. To evaluate detectability, we leverage VideoMAE [tong2022videomae] as the backbone and train it with 100 epochs on a dataset consisting of 8,000 watermark-free videos and 8,000 watermarked videos to perform binary classification. The results in Figure 3, show that the network’s classification accuracy is notably low for videos watermarked with VideoMark, achieving 54.07% on the training set and 48.02% on the validation set. In contrast, other watermarking methods show similar performance on training and validation sets, indicating their watermark patterns are easier to learn and detect.
Figure 3: The binary classification results under different watermarking algorithms.
5.3 Analysis
Impact of Sparsity. As shown in Figure 4, we present the extraction capability of the two models under different sparsity $t$. The extraction accuracy in both tasks reaches its highest value when $t=3$ but drops sharply for other values of $t$. We attribute this phenomenon to an optimal range of $t$, in which decoding is neither affected by redundant signal interference (when $t>3$) nor by insufficient redundancy for correction(when $t<3$). Consequently, we set $t=3$ for all subsequent experiments.
Impact of message length. As shown in Figure 5, we evaluate the extraction capability and robustness across different message lengths. In both tasks, the extraction accuracy remains stable when the message length is below 512, but drops for longer messages, since the redundant bits have reached their limit in providing effective error correction. It indicates that VideoMark can stably embed up to 512 bits of messages and extract them reliably in the absence of attacks. Based on these findings, we fix the message length at 512 bits for all subsequent extraction experiments. Additionally, we evaluate the robustness of VideoMark against three types of spatial tampering. Robustness in the T2V task peaks at 32 watermark bits but degrades as bits increase. Conversely, I2V robustness peaks at 64 bits, slightly surpassing the 32-bit case. We attribute these differences to the higher generative complexity of T2V models, whose greater output variability likely induces more spatial perturbations, making robustness more sensitive to embedding payload.
Figure 4: The binary classification results under different watermarking algorithms.
Figure 5: The extraction accuracy and robustness of VideoMark against spatial tampering for varying message lengths.
Impact of video length. To comprehensively evaluate how the number of generated frames affects watermarking performance, we report both extraction accuracy and visual quality across different generation lengths in Figure 6. We observe that the two metrics degrade to different extents as the number of frames increases from 16 to 32. Extraction accuracy drops from 1.000 to 0.925 in the T2V task and from 0.997 to 0.816 in the I2V task, which indicates that the root cause lies in the increasing difficulty of accurately recovering the original noise as the number of frames rises. This leads to larger cumulative errors and, consequently, a decline in extraction accuracy. Meanwhile, to explore the impact of VideoMark on visual quality, we compare the distribution of video quality scores between watermarked and non-watermarked videos in Figure 7. The distribution of video quality scores for watermarked videos remains consistent with that of clean videos across different frame lengths, which indicates that VideoMark introduces minimal perceptual distortion, regardless of the video length. The primary cause of the video quality degradation is the model’s limited generative capability for longer videos.
Figure 6: Extraction accuracy and video quality scores under varying numbers of generated frames.
Table 5: Bit accuracy with different frame resolution.
[…]
Impact of frame resolution. To evaluate the extraction capability of the watermark at different resolutions, we present the watermark bit accuracy across various resolutions in Table 5. In the T2V task, the extraction accuracy remains at 0.998 even at a resolution of $1280\times 720$, demonstrating strong extraction performance. In contrast, in the I2V task, extraction accuracy peaks at a resolution of $512\times 512$. We attribute this to a balanced trade-off between VideoMark’s error correction capability and inversion errors at this resolution. In contrast, higher resolutions introduce larger inversion errors during the inversion process, which hinder watermark extraction.
Figure 7: Video quality scores under varying numbers of frames, comparing watermarked (w/ wm) and clean (w/o wm) ouputs.
Table 6: Bit accuracy across different inference (Inf.) and inversion (Inv.) steps. The main experimental configuration is marked.
[…]
Impact of the inversion step. To evaluate the impact of mismatch steps between inference and inversion, we present the extraction accuracy at different steps in Table 6. In the T2V task, mismatched steps introduce minimal loss in extraction accuracy, while in the I2V task, they lead to a significant accuracy degradation. We attribute this to the fact that T2V models are typically more robust to small variations in the inversion process, while I2V models are more sensitive to such discrepancies, leading to greater performance degradation. Considering practical implementation efficiency and extraction capability, we fix the inference and inversion steps at 25 for the T2V and 50 for the I2V.
6 Conclusion
In this work, we propose a training-free, undetectable watermarking framework for video diffusion models. Through extensive experiments, we demonstrate that the generated videos retain high visual quality and exhibit no perceptible artifacts attributable to the embedded watermark. However, the current framework relies on approximate inversion techniques, which limit extraction accuracy in certain scenarios. For future improvements, we suggest exploring more advanced or robust inversion algorithms to enhance the reliability and effectiveness of the watermark retrieval process.